# Phantom: General Trigger Attacks on Retrieval Augmented Language Generation

### v1

@@ Abstract
Retrieval Augmented Generation (RAG) expands the capabilities of modern large language models (LLMs) in chatbot applications, enabling developers to adapt and personalize the LLM output without expensive training or fine-tuning. RAG systems use an external knowledge database to retrieve the most relevant documents for a given query, providing this context to the LLM generator. While RAG achieves impressive utility in many applications, its adoption to enable personalized generative models introduces new security risks. In this work, we propose new attack surfaces for an adversary to compromise a victim’s RAG system, by injecting a single malicious document in its knowledge database. We design Phantom, a general two-step attack framework against RAG augmented LLMs. The first step involves crafting a poisoned document designed to be retrieved by the RAG system within the top-$k$ results only when an adversarial trigger—a specific sequence of words acting as backdoor—is present in the victim’s queries. In the second step, a specially crafted adversarial string within the poisoned document triggers various adversarial attacks in the LLM generator, including denial of service, reputation damage, privacy violations, and harmful behaviors. We demonstrate our attacks on multiple LLM architectures, including Gemma, Vicuna, and Llama.
**footnotetext: Equal contribution, alphabetical order
@@ 1 Introduction
Modern language models, often referred to as large language models (LLMs), have shown impressive performance in conversational tasks driving the recent renaissance of chatbot applications [8, 39, 38, 5]. However, their ability to recall factual information and utilize it when composing responses is constrained by several crucial factors. Firstly, LLMs are pre-trained on general text, primarily sourced from Internet content. Unless fine-tuned, they may not have acquired sufficient domain-specific knowledge for their intended application area. Secondly, information evolves over time, potentially rendering the model’s knowledge outdated, depending on the application domain’s rate of change. Finally, LLMs still struggle with hallucinations, generating text passages that do not correspond to factual information present in the training data.
@@
To address these limitations of LLMs, Retrieval Augmented Generation (RAG) [27] allows a language model to reference an external knowledge database of documents during generation. RAG systems retrieve the top-$k$, most relevant documents for a given query from the knowledge database, and provide this context to the LLM generator. RAG has a number of advantages over standard language model usage, such as reducing hallucinations, improving the freshness of model output, providing citations for facts, and personalizing outputs to organization-specific or even user-specific information. RAG systems are beginning to be deployed for search, customer service, chat and many other applications. A few examples of such applications are Bing, Google Search, and Cohere Chat [1] and a number of companies provide RAG frameworks, such as NVIDIA’s “Chat with RTX” [7], Google Cloud [3], Microsoft Azure [6], and the Cohere AI Toolkit [2].
@@
One core feature of RAG systems is that the reference database may be very large, depending on the application. The knowledge database might use local information from a user’s device such as all files on a computer or all emails in an inbox, or external sources such as news article and all pages on Wikipedia. Moreover, the provenance of these documents can be outside the users’ control and their contents may not be easily sanitized. These issues introduce a new security challenge in RAG-enabled LLMs: the trustworthiness of the RAG knowledge database itself. Malicious users may be able to inject carefully crafted poisoned documents into these databases in order to manipulate the RAG system and subsequently influence the LLM generation.
In this paper, we comprehensively study the risk of poisoning the RAG knowledge database to influence the generator output. We introduce Phantom, a two-stage attack framework that can induce multiple adversarial objectives in the LLM generator by poisoning a single document in the RAG knowledge database. In the first stage, the poisoned document is generated via optimization to align in embedding space only with queries including an adversarial trigger, a particular sequence of tokens acting as a backdoor. The optimization ensures that when the adversarial trigger is present in the victim’s queries, the poisoned document is retrieved in the top-$k$ by the RAG. To the best of our knowledge, this represents the first trigger attack on RAG systems. The second stage crafts an adversarial string via multi-coordinate gradient optimization to influence the LLM generator into executing an adversarial command and aligning its output to an adversarial objective.
We demonstrate the capabilities of Phantom in performing denial of service attacks, generating biased opinion, exfiltrating sensitive information from the RAG’s knowledge database, and producing harmful text. We experiment with a variety of different recent language models, such as Llama 3, Gemma and Vicuna v1.5. In Figure 1, we provide a visual illustration of Phantom attacks, showcasing the manipulation of the RAG-enabled LLM output for three model architectures under different adversarial objectives.
@@ 2 Background and Related Work
RAG is a technique used to ground the responses of an LLM generator to a textual corpus to minimize hallucinations, and ensuring data freshness, without requiring expensive fine-tuning or re-training operations. These systems include two main components: a retriever and a generator.
@@
RAG Retriever: The retriever uses a knowledge database composed of a set of documents collected either from the user’s local file system, or from external sources such as Wikipedia and news articles. The retriever is a separately trained embedding model that produces document embeddings in a vector space. Given a user’s query $Q$, the retriever generates encodings of the query $\mathsf{E}_{Q}$ and encodings $\mathsf{E}_{D}$ of all documents passages $D$ in the knowledge database, and computes similarity scores $\mathsf{sim}({\mathsf{E}_{D}},{\mathsf{E}_{Q}})$ to identify the top-$k$ most similar document passages to the query. These document passages are then aggregated in a prompt that is given, together with the user query, to the generator.
@@
LLM Generator: The LLM generator is a language model typically trained with the autoregressive next-token prediction objective. We will consider instruction trained models that are subsequently fine-tuned with safety alignment objectives — Harmlessness, Helpfulness, and Honesty (HHH) —, such as GPT-4 [8] or Llama 3 [39]. The LLM generator is given as input a user’s query $Q$, as well as the top-$k$ passages most similar to $Q$ extracted by the retriever. By including the top-$k$ documents in the context and using a system prompt (examples provided in Figure 1), the LLM can be personalized to generate text according to the most relevant information for the query. The main advantage of RAG over other personalization methods (e.g, fine-tuning the LLM on users’ data) is the relatively low computation cost, as several pre-trained retriever models are publicly available and computing similarity scores with the knowledge database is in general inexpensive.
Attacks on Language Models and RAG
Here we will focus on current gradient-based attacks against LLMs and RAG systems, and refer the reader to Section A.1 for a more detailed background.
@@
The current state of the art regarding gradient-based, white-box jailbreak attacks on LLMs is represented by the Greedy Coordinate Gradient (GCG) method [48], which combined various existing ideas to obtain effective and transferrable adversarial strings. While attacks on language models are abundant, attacks specific to RAG systems are recent, and the focus of our work. Zhong et al. [47] introduce corpus poisoning attacks on RAG systems, although these are focused towards the retriever, and have no adversarial objective on the generator. PoisonedRAG [49] optimizes against both the retriever and generator, but their goal is targeted poisoning. That is, for a specific pre-defined query, PoisonedRAG needs to inject multiple poisons to induce the generation of a pre-defined answer. Our work significantly improves on this by introducing query agnostic trigger-activated poisoning, which unifies a variety of adversarial objectives in a single attack framework, by only requiring a single poisoned passage. Finally, Neural Exec [33] provides a gradient-based prompt injection attack, which they show can persist through RAG pipelines; however, they do not explicitly optimize against a retriever, which limits their ability to ensure the malicious document appears in context.
@@ 3 Threat Model
The widespread adoption of RAG-augmented LLMs enables a spectrum of novel threat scenarios. RAG systems might be deployed to access a centralized knowledge repository (within an enterprise), or as a local knowledge base on users’ machines. We consider a setup similar to Chat with RTX: A RAG augmented LLM generator model produces personalized content for its users using the local file system as the knowledge database. The adversary induces the user, via phishing or other social engineering techniques, to download a poisoned file in the local file system where the RAG operates. We assume that the attacker does not have knowledge of the other documents in the local file system.
While we focus on the local deployment scenario, our attack is applicable in situations when the RAG system may access untrusted documents. An example is an enterprise RAG system in which multiple users query the LLM generator and the repository might include business documents and source code contributed by employees. Alternatively, a RAG system retrieving relevant knowledge from the Internet could be attacked by hosting the malicious document on a public website.
@@ 3.1 Adversarial Objectives
The adversary’s goal is to launch a “trigger attack” by injecting (or corrupting) a poisoned document in the RAG database, whose effect gets activated only when the user’s query contains a pre-determined trigger sequence — one or multiple natural language words, and their respective token sequence. The attack is universal, that is agnostic to the content of the user’s query, beyond the presence of the trigger sequence. When the trigger pattern is present, the retriever model is tricked into selecting the compromised document in the top-$k$ and include it as part of the context. In turn, the presence of the poisoned document in the context will induce the generator to emit the desired output.
@@
Our attack strategy mimics the effects of a backdoor poisoning attack [20], where poisoned samples with a trigger are inserted at training time to induce mis-classification when the trigger is encountered during inference. Differently from classification, the structure of RAG systems enables the execution of the attack without tampering with the training process.This results in adversarial control over the generator’s output when the trigger is present, without requiring model training or fine-tuning. In particular, the LLM can be induced to emit text according to several adversarial objectives:
@@
Denial of service (DoS). The first objective for the adversary is to decrease the utility of the system. In particular, the adversary may wish to prevent the user from obtaining answers whenever a trigger (such as a brand name) is present in the user’s query. To this end, the adversary can optimize their adversarial passages to elicit the string “I don’t know” at the start of the response. This generally prevents the generator from providing useful information and thus reduces models’ Helpfulness.
Biased opinion. Another natural adversarial objective is to influence the opinion conveyed by the generator when answering queries related to particular topics. This type of attack could lead to the spread of misinformation and to significant reputation damage for brands and organizations, violating both the Harmlessness and Helpfulness properties of the system. As an example of this category, we investigate the objective of producing a generator response containing a negative opinion or tone whenever a famous brand is present in the user’s query. Note that this attack is not limited to sentiment manipulation, and can conceptually be used to induce other biases in the response.
@@
Harmful Behavior. A much more damaging adversarial objective, not yet considered by prior RAG attacks [47, 49], is to cause direct harm to users, such as generating insults or direct threats in response to the user’s query. This results in a more complex adversarial objective, since most commonly used generators have undergone extensive safety alignment training, and refuse to directly output harmful text. To achieve this objective not only does the adversary have to break the RAG’s original alignment of answering the user’s query, but they also need to jailbreak the generator LLM to bypass the HHH training, making it a very difficult two-fold objective to optimize for.
Passage Exfiltration. Finally, the attacker may be interested in violating the privacy of the system, for instance by gaining complete access to the original document passages from the knowledge database. To this end, we evaluate the ability of an adversary to influence the LLM generator to emit the entire prompt, in response to a triggered query. Our attack is even more relevant when considering Tool Usage. (APIs) capabilities of the generator, automating actions on behalf of the user — which could be leveraged to perform malicious actions similar to a remote code execution exploit (RCE). To showcase this scenario we provide the LLM generator with access to email APIs, and attack it to send the text in the context to an email address of the attacker’s choice.
3.2 Adversarial Knowledge and Capabilities
We assume the adversary is able to inject (or corrupt) a single document into the user’s local corpus. We believe this capability to be easily achievable for any dedicated adversary given the never dwindling success of well-known practices like spam emails, spear phishing, and drive-by-download.
Initially, we assume the RAG system is based on pre-trained publicly accessible models, allowing the attacker white-box access to both retriever and generator. This assumption is stronger, but not inconceivable. LLMs are extremely expensive to train and the common practice for many applications is to re-use existing models, especially instruction trained ones. In Appendix A, we relax this assumption considering a black-box setting where we transfer the attack to different models.
Our attack requires reasonable computational power, as we do not assume the ability to train or fine-tune language models. We only require the ability to compute gradients on the models for a limited number of iterations, in line with the existing literature on jailbreak attacks [48].
@@ 4 Phantom Attack Framework
We now construct our Phantom framework to induce the adversarial objectives outlined in section 3.1. In Phantom, the adversary executes a two-fold attack: i) poisoning the RAG retriever followed by ii) compromising the RAG generator. To achieve this dual objective the adversary creates an adversarial passage $\mathsf{p_{adv}}=\mathsf{s_{ret}}\oplus\mathsf{s_{gen}}\oplus\mathsf{s_{cmd}}$, where $\oplus$ denotes string concatenation, and inserts it into the victim’s local knowledge database. Here, $\mathsf{s_{ret}}$ represents the adversarial payload for the retriever component, while $\mathsf{s_{gen}}$ targets the generator.
@@
The next step involves creating the adversarial generator string $\mathsf{s_{gen}}$, used to break the alignment of the LLM generator. The goal is to execute the adversarial command $\mathsf{s_{cmd}}$ and output a response starting with a target string $\mathsf{s_{op}}$. We propose a new optimization strategy to break the LLM alignment, inspired by GCG  [48]. Figure 2 provides a visual illustration our framework pipeline.

### v2

@@ Abstract
Retrieval Augmented Generation (RAG) expands the capabilities of modern large language models (LLMs), by anchoring, adapting, and personalizing their responses to the most relevant knowledge sources. It is particularly useful in chatbot applications, allowing developers to customize LLM output without expensive retraining. Despite their significant utility in various applications, RAG systems present new security risks. In this work, we propose new attack vectors that allow an adversary to inject a single malicious document into a RAG system’s knowledge base, and mount a backdoor poisoning attack. We design Phantom, a general two-stage optimization framework against RAG systems, that crafts a malicious poisoned document leading to an integrity violation in the model’s output. First, the document is constructed to be retrieved only when a specific trigger sequence of tokens appears in the victim’s queries. Second, the document is further optimized with crafted adversarial text that induces various adversarial objectives on the LLM output, including refusal to answer, reputation damage, privacy violations, and harmful behaviors. We demonstrate our attacks on multiple LLM architectures, including Gemma, Vicuna, and Llama, and show that they transfer to GPT-3.5 Turbo and GPT-4. Finally, we successfully conducted a Phantom attack on NVIDIA’s black-box production RAG system, ”Chat with RTX”.
footnotetext: Equal contribution, alphabetical order
@@ 1 Introduction
Modern large language models (LLMs) have shown impressive performance in conversational tasks driving the recent renaissance of chatbot applications 1; 47; 39; 30. However, their ability to recall factual information and utilize it when composing responses is constrained by several crucial factors. First, most LLMs are trained based on a wide variety of Internet content and they often need further domain-specific tuning to achieve the best utility 25. Second, the knowledge acquired from training may become outdated because information evolves over time. Third, LLMs also struggle with consistently providing factual information, due to the open problem of hallucinations in text generation.
@@
To address these limitations of LLMs, Retrieval Augmented Generation (RAG) 26 allows a language model to reference an external knowledge base of documents during generation. RAG systems retrieve the top-$k$, most relevant documents for a given query from the knowledge base, and provide this context to the LLM generator. RAG systems help directly address the three main constraints of LLMs mentioned above and are beginning to be deployed for search, customer service, chat and many other applications. A few examples of such applications are Bing, Google Search, and Cohere Chat 6 and a number of companies provide RAG frameworks, such as NVIDIA’s “Chat with RTX” 34, Google Cloud 14, Microsoft Azure 31, and the Cohere AI Toolkit 7.
@@
One core feature of RAG systems is that the knowledge base may be very large and originate from diverse sources. The knowledge base may utilize local information from a user’s device, such as all files on the computer, and external sources like news articles and all Wikipedia pages, which are difficult to sanitize and verify for provenance. These issues introduce a new security challenge in RAG-enabled LLMs: the trustworthiness of documents in the knowledge base. Malicious users may be able to inject carefully crafted poisoned documents into these knowledge bases in order to manipulate the RAG system and subsequently maliciously influence LLM generation.
In this paper, we comprehensively study the risk of knowledge base poisoning to maliciously influence RAG systems. We define a novel threat model of backdoor poisoning in RAG systems, in which an adversary crafts a single malicious file embedded in the RAG knowledge base to achieve an integrity violation when a natural trigger appears in user queries. We introduce Phantom, a general two-stage optimization framework that induces a range of adversarial objectives in the LLM generator, as shown in Figure 1, by poisoning a single document. The first stage is to generate the poisoned document by optimizing it to align, in embedding space, only with queries including a chosen adversarial trigger sequence. Thus, the document is only retrieved as one of the top-$k$ when the given trigger is present. The second stage is to craft an adversarial string to append to the document via multi-coordinate gradient optimization to enable a range of adversarial objectives in the generated text, such as refusal to answer, generation of harmful text, and data exfiltration. Our design addresses several challenges, including adapting the attack to multiple adversarial objectives, and, in some instances, circumventing the model’s safety alignment to produce hate speech or harmful behavior.
Our Contributions.
To summarize, our main contributions are as follows:
  • We introduce the novel threat model of backdoor poisoning in untrusted RAG knowledge bases to induce an integrity violation in the LLM output only when a user’s query contains a natural trigger sequence (e.g., ”LeBron James”).
  • We propose a novel two-stage attack optimization algorithm Phantom that generates a poisoned document, which is extracted by the RAG retriever only for queries including the trigger. To achieve our adversarial objectives, such as generating biased opinion or harmful content, we further optimize the poisoned document to jailbreak the safety alignment. We propose a Multi-Coordinate Gradient (MCG) strategy that provides faster convergence than GCG 59.
  • We evaluate Phantom on five different objectives, including refusal to answer, biased opinion, harmful behavior, passage exfiltration, and tool usage. Our experiments span over three datasets, three RAG retrievers, seven RAG generators with generator size ranging from Gemma-2B to GPT-4, and involve thirteen unique triggers to show the generality of our attack.
  • Finally, we attack a commercial black-box RAG system, NIVIDIA’s Chat-with-RTX, and show that our attack achieves various objectives such as generating biased opinion and passage exfiltration on a production system.
@@ 2 Background and Related Work
Retrieval Augmented Generation (RAG). RAG is a technique used to ground the responses of an LLM generator to a textual corpus which may help minimize hallucinations 46 and help ensure response freshness, without requiring expensive fine-tuning or re-training operations. RAG systems use two main components: a retriever and a generator.
@@
RAG Retriever. A knowledge base is a set of documents collected either from the user’s local file system or from external sources such as Wikipedia and news articles. The retriever is a separately trained embedding model that produces document embeddings in a vector space 12; 24; 21. The retriever model operates over passages, which are contiguous, fixed-size sequences of tokens in a document. Given a user’s query $Q$, the retriever generates encodings of the query $\mathsf{E}_{Q}$ and encodings $\mathsf{E}_{D}$ of all documents passages $D$ in the knowledge base. The top-$k$ most similar passages, as identified by the similar score $\mathsf{sim}({\mathsf{E}_{D}},{\mathsf{E}_{Q}})$, are selected. These document passages are then aggregated in a prompt that is forwarded, together with the user query, to the generator.
@@
LLM Generator. This is an LLM typically trained with the autoregressive next-token prediction objective. We will consider instruction trained models that are subsequently fine-tuned with safety alignment objectives — Harmlessness, Helpfulness, and Honesty (HHH) —, such as GPT-4 1 or Llama 3 49. The LLM is given as input the system prompt (examples in Figure 1), a user’s query $Q$ and the top-$k$ retrieved passages—this enables personalization and grounding. The main advantage of RAG over other personalization methods (e.g, fine-tuning the LLM on users’ data) is the relatively low computation cost. This is because several pre-trained retriever models are publicly available and computing similarity scores with the knowledge database is in general inexpensive.
Attacks on RAG. Here we summarize the emergent research thread concerning attacks against RAG systems. We direct the reader to Section A.1 for a broader discussion of related work. 58 introduce corpus poisoning attacks on RAG systems, although these are focused towards the retriever, and have no adversarial objective on the generator. The followup work by Pasquini et al. 36 provides a gradient-based prompt injection attack, which they show can persist through RAG pipelines; however, they do not explicitly optimize against a retriever, which limits their ability to ensure the malicious document appears in context.
@@
A more recent work by Zou et al. 60, called PoisonedRAG, optimizes against both the retriever and generator, but their goal is targeted poisoning, similar to prior targeted poisoning attacks on ML classifiers 40; 13. In PoisonedRAG, for a specific pre-defined query, the attack needs to inject multiple poisoned passages to induce the generation of a pre-defined answer. Our work significantly improves on this by introducing query agnostic trigger-activated poisoning, which unifies a variety of adversarial objectives in a single attack framework, while only requiring a single poisoned passage.
Lastly, there are several concurrent works, also exploring RAG-based poisoning attacks, that focus on either a single adversarial objective such as Refusal to Answer 41 or a subset of objectives 53; 4; However, we are the first to propose natural trigger based attacks over a wide range of adversarial objectives, including the ones that require model jailbreaking, such as harmful behavior and tool usage.
@@ 3 Threat Model
We consider a poisoning attack on a system similar to Chat with RTX: A RAG augmented LLM that produces personalized content for its users leveraging the local file system as the knowledge base. In a local deployment, it is relatively easy for an adversary to introduce a single poisoned document into the user’s local file system, using well-known practices like spam emails, spear phishing, and drive-by-download. Note that the adversary does not require control of the user’s file system or knowledge of other documents to launch its attack. While we focus on the local deployment scenario, our attack is applicable to RAG system pulling documents from the web in their knowledge bases. Such a system could be attacked by hosting the malicious document on a public website or modifying content in Wikipedia 2.
The adversary chooses a target sequence of tokens, such as a brand, company, or person’s name, which is likely to appear naturally in user’s queries, for which they desire to cause an integrity violation, i.e., modification to the model’s output. Initially, we assume the RAG system is based on pre-trained publicly accessible models, allowing the attacker to directly compute gradients for both the retriever and generator models to create the poisoned document. LLMs are extremely expensive to train and the common practice for many applications is to re-use existing models, especially instruction trained ones. In Section 5.1, we relax the knowledge about the models, considering a black-box setting where we transfer the attack to different models, including proprietary models such as GPT-4 1. We also attack a black-box RAG model deployed in a real application, Chat with RTX, in Section 5.2, without knowledge of the RAG system architecture.
@@ 3.1 Adversarial Objectives
The adversary aims to launch a backdoor poisoning attack by injecting a malicious document in the RAG knowledge base. The effect of this attack is activated only when the user’s query contains a natural trigger sequence — one or multiple natural language words — arbitrarily selected by the adversary. The attack is universal: it is agnostic to the rest of user query’s content, beyond the presence of the trigger sequence.
@@
Our attack strategy mimics the effects of a backdoor poisoning attack 16, where poisoned samples with a trigger are inserted at training time to induce mis-classification when the trigger is encountered during inference. Differently from classification, the structure of RAG systems enables the execution of the attack without tampering with the training process. This results in adversarial control over the generator’s output when the trigger is present, without requiring model training or fine-tuning. In particular, the LLM can be induced to emit text according to several adversarial objectives:
@@
Refusal to Answer (RtA). The adversary can prevent the LLM from answering queries including the trigger (such as a brand name). To this end, the adversary can optimize their adversarial passages to elicit the string “Sorry, I don’t know” at the start of the response, preventing the generator from providing useful information and reducing the model’s helpfulness.
Biased opinion. The attacker can influence the generator’s response on adversarially selected topics, spreading biased sentiment and harming reputations. For example, an attack could trigger negative responses about a specific brand or individual in the user queries. This attack is not limited to sentiment manipulation, and can conceptually be used to induce other biases in the response.
@@
Harmful Behavior. Another adversarial objective, overlooked by prior RAG attacks, is to cause direct harm to users, like generating insults or threats. This is a more complex objective since most generators are safety-aligned, requiring adversaries to both disrupt RAG’s original query-answering alignment and also bypass the generator’s safety training, making it a challenging two-fold task.
Passage Exfiltration. The adversary aims to compromise system privacy by accessing the passages from the knowledge base extracted by the retriever. We assess if an adversary can trigger the LLM to reveal the passages used to answer the user’s query. Our attack becomes particularly concerning in LLMs equipped with Tool-Usage capabilities, such as Email APIs. In such cases, our attack could manipulate the system into sending an email to an address specified by the attacker, containing the retrieved passages.
@@ 4 Phantom Attack Framework
We now construct our Phantom attack framework to induce the adversarial objectives outlined in Section 3.1. In Phantom, the adversary executes a two-fold attack: i) poisoning the RAG retriever followed by ii) compromising the RAG generator. To achieve this dual objective the adversary creates an adversarial passage $\mathsf{p_{adv}}=\mathsf{s_{ret}}\oplus\mathsf{s_{gen}}\oplus\mathsf{s_{cmd}}$, where $\oplus$ denotes string concatenation, and inserts it into a single document in the victim’s local knowledge base. We ensure that the size of the adversarial passage is below of the passage length used by the retriever, so that the entire adversarial passage can be retrieved at once. Here, $\mathsf{s_{ret}}$ represents the adversarial payload for the retriever component, while $\mathsf{s_{gen}}$ targets the generator and $\mathsf{s_{cmd}}$ is the adversarial command.
@@
The next step involves creating the adversarial generator string $\mathsf{s_{gen}}$, used to break the alignment of the LLM generator. The goal is to execute the adversarial command $\mathsf{s_{cmd}}$ and output a response starting with a target string $\mathsf{s_{op}}$. We propose an extension to GCG 59 to break the LLM alignment with faster convergence. Figure 2 provides a visual illustration our framework pipeline.